Danny “Dimes” Donahue is a neighborhood’s 9-year-old entrepreneur. His most recent venture is selling homemade brownies that he bakes himself. At a price of $1.5 each, he sells 100. At a price of $1 each, he sells 300. Is demand elastic or inelastic over this price range? If demand had the same elasticity for a price decline from $1 to $0.5 as it does for the decline from $1.5 to $1, would cutting the price from $1 to $0.5 increase or decrease Danny’s total revenue?


1
Expert's answer
2020-10-21T11:09:26-0400

(i) Calculation to show whether demand is elastic or inelastic.

Elasticity of demand, "E_{d}" = %change in quantity demanded "\div" %change in price

%change in demand = "\frac{Q_{2} - Q_{1}}{Q_{1}}\times100= \frac{300-100}{100}\times100=200"

%change in price= "\frac{P_{2}-P_{1}}{P_{1}}\times100= \frac{1-1.5}{1.5}\times100=-33.33"

"E_{d}= -\frac{200}{33.33}= -6"

The demand is therefore elastic since the coefficient of price elasticity of demand is greater than one.

(ii) when the demand is elastic, the relationship between total revenue and price is always such that, a fall in price lead to an increase in revenue. therefore, the total revenue will increase if the price is cut from $1 to $0.5.